# Traceable Black-Box Watermarks For Federated Learning

### v3

@@ Abstract
Due to the distributed nature of Federated Learning (FL) systems, each local client has access to the global model, posing a critical risk of model leakage. Existing works have explored injecting watermarks into local models to enable intellectual property protection. However, these methods either focus on non-traceable watermarks or traceable but white-box watermarks. We identify a gap in the literature regarding the formal definition of traceable black-box watermarking and the formulation of the problem of injecting such watermarks into FL systems. In this work, we first formalize the problem of injecting traceable black-box watermarks into FL. Based on the problem, we propose a novel server-side watermarking method, $\mathbf{TraMark}$, which creates a traceable watermarked model for each client, enabling verification of model leakage in black-box settings. To achieve this, $\mathbf{TraMark}$ partitions the model parameter space into two distinct regions: the main task region and the watermarking region. Subsequently, a personalized global model is constructed for each client by aggregating only the main task region while preserving the watermarking region. Each model then learns a unique watermark exclusively within the watermarking region using a distinct watermark dataset before being sent back to the local client. Extensive results across various FL systems demonstrate that $\mathbf{TraMark}$ ensures the traceability of all watermarked models while preserving their main task performance.
@@ 1 Introduction
Federated Learning (FL) is a promising training paradigm that enables collaborative model training across distributed local clients while ensuring that private data remains on local devices (McMahan et al. 2017). Instead of sharing raw data, clients train local models independently and periodically send updates to a central server, which aggregates them into a global model. This privacy-preserving benefit has led to FL’s widespread adoption in various fields, including healthcare (Nguyen et al. 2022), finance (Long et al. 2020), and remote sensing (Liu et al. 2020), where local data privacy is a critical concern. However, sharing the global model with all participants introduces risks of model leakage. Specifically, malicious clients may exploit their access by duplicating and illegally distributing the model (Li et al. 2022). Such misconduct undermines the integrity of the FL system and compromises the collective interests of all participants. Consequently, protecting intellectual property (IP) (Xue et al. 2022) rights of FL-trained models and detecting copyright infringement have become critical challenges in FL.
@@
Protecting the IP of FL-trained models requires mechanisms to verify rightful ownership if a model is unlawfully distributed (Shao et al. 2024) (i.e., proving that the model originated from the FL system). To address this, researchers have proposed embedding watermarks into the global model to enable ownership verification. Existing approaches primarily fall into two categories: parameter-based (Uchida et al. 2017) and backdoor-based (Tekgul et al. 2021; Adi et al. 2018) watermarking techniques. Parameter-based methods embed signatures (e.g., bit strings) within the model’s parameters as a secret key. During verification, the verifier extracts this key and applies a cryptographic function with the corresponding public key to validate the model’s ownership. However, this process requires white-box access to the model parameters, which is often impractical, particularly when the suspect model is only accessible in a black-box setting (e.g., via an API). To overcome this limitation, backdoor-based watermarking leverages backdoor injection techniques to ensure that the model learns a specific trigger. A watermarked model outputs predefined responses when presented with inputs containing the trigger. Unlike parameter-based methods, this verification process does not require access to the model parameters, making backdoor-based watermarking a more practical solution.
@@
Beyond ownership verification, the verifier also needs to trace the source of model leakage, identifying which client was responsible for the unauthorized distribution. Recent studies have explored methods for ensuring the traceability of watermarked models in FL (Shao et al. 2024; Yu et al. 2023; Xu et al. 2024; Nie and Lu 2024a). For instance, FedTracker (Shao et al. 2024) embeds unique bit strings into each client’s model and identifies the leaker by measuring bit string similarities. However, this approach requires white-box access to the suspect model, which limits its practicality. Another approach, FedCRMW (Nie and Lu 2024a), introduces a black-box watermarking mechanism that injects unique watermarks into each model shared with the client by mixing clients’ local datasets with multiple types of triggers. The model leaker is then identified based on the models’ predictions on these watermarked datasets. However, this method modifies the local training protocol and requires access to clients’ local data, making it vulnerable to tampering by malicious clients. Moreover, such an approach contradicts FL’s core principle of data privacy preservation. Despite these advancements, the literature lacks a formal definition and formulation of the problem of black-box watermarking for ownership verification and traceability in FL.
@@
In this work, we first formalize the traceable black-box watermarking injection problem in an FL system. Building on this, we propose $\mathbf{TraMark}$ , which creates a personalized, traceable watermarked model for each client, enabling verification of model leakage in black-box settings. Specifically, to ensure traceability, the server partitions the model’s parameter space into two regions: the main task region, responsible for learning the primary FL task, and the watermarking region, designated for embedding a watermark. The server then generates personalized global models for each client via masked aggregation. Each model is subsequently injected with a distinct watermark exclusively in the watermarking region using a dedicated watermark dataset. This process ensures that every client receives a personalized global model that integrates aggregated knowledge from other clients while embedding a unique watermark for model leakage verification. We summarize the contribution of our paper as follows.
@@
  • To the best of our knowledge, this is the first work to formally formulate the problem of traceable black-box watermark injection in FL systems. Based on this, we propose $\mathbf{TraMark}$, a novel watermarking method that seamlessly integrates into existing FL systems.
@@ 2 Background and System Settings
We provide a notation table in Table 2 in Appendix A.1.
Federated Learning.
A typical FL system consists of a central server and a set of $n$ local clients, which collaboratively train a shared model $\theta\in\mathbb{R}^{d}$. The FL problem is generally formulated as: $\min_{\theta}(1/n)\sum_{i=1}^{n}F_{i}(\theta;\mathcal{D}^{l}_{i}),$ where $F_{i}(\cdot)$ represents the local learning objective of client $i$, and $\mathcal{D}^{l}_{i}$ is its local dataset. For instance, for a classification task, client $i$’s local objective can be expressed as: $F_{i}(\theta;\mathcal{D}^{l}_{i})\coloneq\mathbb{E}_{(z,y)\in\mathcal{D}^{l}_{i}}\mathcal{L}(\theta;z,y),$ where $\mathcal{L}(\cdot)$ is the loss function, and $(z,y)$ represents a datapoint sampled from $\mathcal{D}^{l}_{i}$. A classic method to solve the FL problem is Federated Averaging (FedAvg) (McMahan et al. 2017). Specifically, in each training round $t$, the server broadcasts the current global model $\theta^{t}$ to each client $i\in[n]$. Upon receiving $\theta^{t}$, client $i$ performs $\tau_{l}$ iterations of local training on it using $\mathcal{D}^{l}_{i}$, resulting in an updated local model $\theta^{t,\tau}_{i}$. The client then computes and sends the model update $\Delta^{t}_{i}=\theta^{t,\tau}_{i}-\theta^{t}$ back to the server. The server aggregates these updates from all clients and refines the global model as: $\theta^{t+1}=\theta^{t}+(1/n)\sum_{i=1}^{n}\Delta^{t}_{i}.$ This process repeats until the global model converges.
Related Work.
We provide a detailed related work and the comparison of our proposed method $\mathbf{TraMark}$ with existing methods in Appendix A.2.
Attack Model.
In this work, we consider a widely used attack model where all clients in an FL system are potential model leakers (Shao et al. 2024; Tekgul et al. 2021). Specifically, we assume these clients will follow a predefined local training protocol to complete the FL task. However, they may illegally distribute their local models for personal profit.
Defense Model.
We assume that the server acts as the defender, responsible for injecting traceable watermarks to the FL system. The server is considered always reliable and equipped with sufficient computational resources. It is also assumed to have full access to all local models but no access to local data. Moreover, the server aims to keep the watermark injection process confidential from all local clients. Additionally, the server also acts as the verifier: if a model is deemed suspicious, it initiates a verification process to determine whether the model originates from the FL system and to identify the responsible model leaker.
@@ 3 Problem Formulation
Black-box Watermarking.
A black-box watermark is a practical watermarking solution that the verification process that does not require access to model parameters, making it more suitable for real-world deployment than white-box watermarking. A common black-box watermarking approach leverages backdoor injection (Adi et al. 2018), where a watermark dataset $\mathcal{D}^{w}$ (as defined in Definition 1) containing triggers is used to train the model to produce a predefined output when presented with the triggers.
@@
where $\phi(x)$ is the unique predefined output distribution assigned to each trigger $x$ in the trigger set $\mathcal{X}^{w}$.
@@
where $\mathbf{y}(\theta^{\prime};x)$ denotes the output probability distribution of the watermarked model given trigger $x$.
If a black-box watermark is successfully embedded, one can verify whether a suspicious model originates from the system by testing its outputs on triggers from $\mathcal{D}^{w}$. For example, in classification tasks, verification is typically performed by evaluating the prediction accuracy of the suspicious model $\theta^{\prime}$ on $\mathcal{D}^{w}$. Specifically, if the model’s prediction accuracy $\sum_{x\in\mathcal{D}^{w}}\mathbf{1}[\arg\max\mathbf{y}(\theta^{\prime};x)=\arg\max\phi(x)]/|\mathcal{D}^{w}|$ exceeds a predefined threshold $\nu$, this indicates that the suspicious model contains the watermark $\delta$, thereby verifying its ownership (Tekgul et al. 2021; Shao et al. 2024; Liu et al. 2021; Nie and Lu 2024b).
Traceability.
However, ownership verification alone is insufficient for detecting model leaking of an FL system. While ownership verification confirms whether a model originates from the system, it does not identify which client leaked it. The ability to pinpoint the source of leakage is known as traceability. To achieve traceability, each watermarked model should carry a distinct watermark, ensuring that every client receives a unique identifier embedded in their model. More formally, for any two watermarked models $\theta^{\prime}_{i}$ and $\theta^{\prime}_{j}$, their outputs should be as different as possible when evaluated on the same watermark dataset. If their outputs are too similar, a watermark collision (as defined in Definition 3) occurs, which can compromise traceability.
@@
$\mathbb{E}_{x}\big{[}\mathrm{Div}\big{(}\mathbf{y}(\theta^{\prime}_{i};x),\mathbf{y}(\theta^{\prime}_{j};x)\big{)}\big{]}\leq\sigma,$
for $x\in\mathcal{D}^{w}_{i}$ or $x\in\mathcal{D}^{w}_{j}$, where $\mathrm{Div}(\cdot)$ is a divergence measurement function (e.g., KL divergence), and $\sigma$ is a predefined collision threshold.
@@
$\mathbb{E}_{x\in\mathcal{D}^{w}_{i}}\big{[}\mathrm{Div}\big{(}\mathbf{y}(\theta^{\prime}_{i};x),\mathbf{y}(\theta^{\prime}_{j};x)\big{)}\big{]}>\sigma,\quad\forall j\in[n],j\neq i.$
@@
Problem Formulation.
Now, we define the problem of injecting traceable black-box watermarks in FL in Problem 1.
@@
$s.t.\ \mathbb{E}_{x\in\mathcal{D}^{w}_{i}}[\mathrm{Div}(\mathbf{y}(\theta+\delta_{i};x),\mathbf{y}(\theta+\delta_{j};x))]>\sigma,$
$\quad\quad\quad\quad\ \ \ \quad\quad\quad\quad\quad\quad\quad\quad\forall i,j\in[n],\ i\neq j,$
@@
From Problem 1, we derive the following key insights: 1) A straightforward solution for Problem 1 is to offload each watermark dataset $\mathcal{D}^{w}_{i}$ to client $i$, allowing clients to mix $\mathcal{D}^{w}_{i}$ with their main task dataset $\mathcal{D}^{l}_{i}$ during local training to solve both objectives simultaneously, as proposed in (Nie and Lu 2024a, b; Liu et al. 2021; Xu et al. 2024; Wu et al. 2022; Li et al. 2022). However, this method is highly vulnerable to malicious clients who may simply discard $\mathcal{D}^{w}_{i}$, leading to the absence of watermarks in their models. Furthermore, if malicious clients are aware of the watermarking process, they could intentionally tamper with it, undermining its effectiveness. To mitigate these risks, it is preferable to decompose Problem 1, leaving the main task to local clients while performing watermark injection solely on the server. 2) A critical challenge in watermark injection is the risk of watermark collisions due to model averaging during aggregation. Specifically, even if the server successfully injects distinct watermarks into the models before sending them to clients for local training, these watermarks will be fused during model aggregation in the next training round if parameters from all clients are simply averaged, as in FedAvg. To address this issue, a specialized mechanism is required to prevent watermark entanglement during aggregation. 3) The constraint in Problem 1 suggests that to avoid collisions, the server should maximize $\|\delta_{i}-\delta_{j}\|^{2}_{2}$. However, if $\delta_{i}$ and $\delta_{j}$ become too divergent, this may impair the main task performance of the watermarked models. Additionally, since the server lacks access to clients’ local datasets, directly solving the watermarking objective $L(\cdot)$ could also lead to significant degradation in the performance of the main task. Thus, a specialized learning strategy is required to ensure that watermark injection does not compromise the model’s main task performance.

### v4

@@ Abstract
Due to the distributed nature of Federated Learning (FL) systems, each local client has access to the global model, which poses a critical risk of model leakage. Existing works have explored injecting watermarks into local models to enable intellectual property protection. However, these methods either focus on non-traceable watermarks or traceable but white-box watermarks. We identify a gap in the literature regarding the formal definition of traceable black-box watermarking and the formulation of the problem of injecting such watermarks into FL systems. In this work, we first formalize the problem of injecting traceable black-box watermarks into FL. Based on the problem, we propose a novel server-side watermarking method, $\mathbf{TraMark}$, which creates a traceable watermarked model for each client, enabling verification of model leakage in black-box settings. To achieve this, $\mathbf{TraMark}$ partitions the model parameter space into two distinct regions: the main task region and the watermarking region. Subsequently, a personalized global model is constructed for each client by aggregating only the main task region while preserving the watermarking region. Each model then learns a unique watermark exclusively within the watermarking region using a distinct watermark dataset before being sent back to the local client. Extensive results across various FL systems demonstrate that $\mathbf{TraMark}$ ensures the traceability of all watermarked models while preserving their main task performance. The code is available at https://github.com/JiiahaoXU/TraMark.
@@ 1 Introduction
Federated Learning (FL) is a promising training paradigm that enables collaborative model training across distributed local clients while ensuring that private data remains on local devices (McMahan et al., 2017). Instead of sharing raw data, clients train local models independently and periodically send updates to a central server, which aggregates them into a global model. This privacy-preserving benefit has led to FL’s widespread adoption in various fields, including healthcare (Nguyen et al., 2022), finance (Long et al., 2020), and remote sensing (Liu et al., 2020), where local data privacy is a critical concern. However, sharing the global model with all participants introduces risks of model leakage. Specifically, malicious clients may exploit their access by duplicating and illegally distributing the model (Li et al., 2022a). Such misconduct undermines the integrity of the FL system and compromises the collective interests of all participants. Consequently, protecting intellectual property (IP) rights of FL-trained models and detecting copyright infringement have become critical challenges in FL (Xue et al., 2022).
@@
Protecting the IP of FL-trained models requires mechanisms to verify rightful ownership if a model is unlawfully distributed (Shao et al., 2024) (i.e., proving that the model originated from the FL system). To address this, researchers have proposed embedding watermarks into the global model to enable ownership verification. Existing approaches primarily fall into two categories: parameter-based (Uchida et al., 2017) and backdoor-based (Tekgul et al., 2021; Adi et al., 2018) watermarking techniques. Parameter-based methods embed signatures (e.g., bit strings) within the model’s parameters as a secret key. During verification, the verifier extracts this key from the suspect model and applies a cryptographic function with the corresponding public key to validate the model’s ownership. However, this process requires white-box access to the model parameters, which is often infeasible in practice, particularly when the suspect model is only accessible in a black-box setting (e.g., via an API). To overcome this limitation, backdoor-based watermarking leverages backdoor injection techniques to ensure that the model learns a specific trigger. A watermarked model outputs predefined responses when presented with inputs containing the trigger. Unlike parameter-based methods, this verification process does not require access to the model parameters, making backdoor-based watermarking a more practical solution.
@@
Beyond ownership verification, the verifier also needs to trace the source of model leakage, identifying which client was responsible for the unauthorized distribution. Recent studies have explored methods for ensuring the traceability of watermarked models in FL (Shao et al., 2024; Yu et al., 2023; Xu et al., 2024; Nie and Lu, 2024a). For instance, FedTracker (Shao et al., 2024) embeds unique bit strings into each client’s model and identifies the leaker by measuring bit string similarities. However, this approach requires white-box access to the suspect model, which limits its practicality. Another approach, FedCRMW (Nie and Lu, 2024a), introduces a black-box watermarking mechanism that injects unique watermarks into each model shared with the client by mixing clients’ local datasets with multiple types of triggers. The model leaker is then identified based on the models’ predictions on these watermarked datasets. However, this method modifies the local training protocol and requires access to clients’ local data, making it vulnerable to tampering by malicious clients. A recent method, MFL-Owner (Gai et al., 2025), targets multi-modal FL by leveraging each client’s visual and language encoders to construct an orthogonal transformation on the client’s trigger set, which serves as a watermark. However, this approach suffers from poor generalizability and fails to scale effectively to broader FL systems. Despite notable empirical progress, existing literature still lacks a formal problem definition and formulation of black-box watermarking for ownership verification and traceability in FL.
@@
In this work, we first formalize the problem of traceable black-box watermarking injection in FL systems. Building on this, we propose $\mathbf{TraMark}$, which creates a personalized, traceable watermarked model for each client, enabling verification of model leakage in black-box settings. Specifically, to ensure traceability, the server partitions the model’s parameter space into two regions: the main task region, responsible for learning the primary FL task, and the watermarking region, designated for embedding a watermark. The server then generates personalized global models for each client via masked aggregation. Each model is subsequently injected with a distinct watermark exclusively in the watermarking region using a dedicated watermark dataset. This process ensures that every client receives a personalized global model that integrates aggregated knowledge from other clients while embedding a unique watermark for model leakage verification. We summarize the contributions of our paper as follows.
@@
  • To the best of our knowledge, this is the first work to formally formulate the problem of traceable black-box watermarking in FL systems. Based on this, we propose $\mathbf{TraMark}$, a novel watermarking method that can be seamlessly integrated into existing FL systems.
@@ 2 Background and System Settings
Federated Learning. A typical FL system consists of a central server and a set of $n$ local clients, which collaboratively train a shared model $\theta\in\mathbb{R}^{d}$. The FL problem is generally formulated as: $\min_{\theta}(1/n)\sum_{i=1}^{n}F_{i}(\theta;\mathcal{D}^{l}_{i}),$ where $F_{i}(\cdot)$ represents the local learning objective of client $i$, and $\mathcal{D}^{l}_{i}$ is its local dataset. For instance, for a classification task, client $i$’s local objective can be expressed as: $F_{i}(\theta;\mathcal{D}^{l}_{i})\coloneq\mathbb{E}_{(z,y)\in\mathcal{D}^{l}_{i}}\mathcal{L}(\theta;z,y),$ where $\mathcal{L}(\cdot)$ is the loss function, and $(z,y)$ represents a datapoint sampled from $\mathcal{D}^{l}_{i}$. A classic method to solve the FL problem is Federated Averaging (FedAvg) (McMahan et al., 2017). Specifically, in each training round $t$, the server broadcasts the current global model $\theta^{t}$ to each client $i\in[n]$. Upon receiving $\theta^{t}$, client $i$ performs $\tau_{l}$ iterations of local training on it using $\mathcal{D}^{l}_{i}$, resulting in an updated local model $\theta^{t,\tau}_{i}$. The client then computes and sends the model update $\Delta^{t}_{i}=\theta^{t,\tau}_{i}-\theta^{t}$ back to the server. The server aggregates updates from all clients and refines the global model as: $\theta^{t+1}=\theta^{t}+(1/n)\sum_{i=1}^{n}\Delta^{t}_{i}.$ This process repeats until the global model converges.
Attack Model. In this work, we consider a widely used attack model where all clients in an FL system are potential model leakers (Shao et al., 2024; Tekgul et al., 2021). Specifically, we assume these clients will follow a predefined local training protocol to complete the FL task. However, they may illegally distribute their local models for personal profit.
Defense Model. In this work, we consider the standard FL setup (e.g., FedAvg) where the server directly receives and aggregates local model updates but has no access to any private local data. The server is considered always reliable and equipped with sufficient computational resources. We assume that the server acts as the defender, responsible for injecting traceable watermarks into the FL system. Moreover, the server aims to keep the watermark injection process confidential from all local clients. Additionally, the server also acts as the verifier: if a model is deemed suspicious, it initiates a verification process to determine whether the model originates from the FL system and to identify the responsible model leaker.
@@ 3 Problem Formulation
Black-box Watermarking. A black-box watermark is a practical watermarking solution that enables verification without access to model parameters, making it more suitable for real-world deployment. A common black-box watermarking approach leverages backdoor injection (Adi et al., 2018), where a watermark dataset $\mathcal{D}^{w}$ (as defined in Definition 1) containing triggers is used to train the model to produce a predefined output when presented with the triggers.
@@
where $\phi(x)$ is the unique predefined output distribution assigned to each trigger $x$ in trigger set $\mathcal{X}^{w}$.
@@
where $\mathbf{y}(\theta^{\prime};x)$ is the output probability distribution of the watermarked model given a trigger $x$.
If a black-box watermark is successfully embedded, one can verify whether a suspicious model originates from the system by testing its outputs on triggers from $\mathcal{D}^{w}$. For example, in classification tasks, verification is typically performed by evaluating the prediction accuracy of the suspicious model $\theta^{\prime}$ on $\mathcal{D}^{w}$. Specifically, if the model’s prediction accuracy $\sum_{x\in\mathcal{D}^{w}}\mathbf{1}[\arg\max\mathbf{y}(\theta^{\prime};x)=\arg\max\phi(x)]/|\mathcal{D}^{w}|$ exceeds a predefined threshold $\nu$, this indicates that the suspicious model contains the watermark $\delta$, thereby verifying its ownership (Tekgul et al., 2021; Shao et al., 2024; Liu et al., 2021; Nie and Lu, 2024b).
Traceability. However, ownership verification alone is insufficient for detecting model leaking of an FL system. While ownership verification confirms whether a model originates from the system, it does not identify which client leaked it. The ability to pinpoint the source of leakage is known as traceability. To achieve traceability, each watermarked model should carry a distinct watermark, ensuring that every client receives a unique identifier embedded in their model. More formally, for any two watermarked models $\theta^{\prime}_{i}$ and $\theta^{\prime}_{j}$, their outputs should be as different as possible when evaluated on the same watermark dataset. If their outputs are too similar, a watermark collision (as defined in Definition 3) occurs, which can compromise traceability.
@@
$\mathbb{E}_{x}\big[\mathtt{Div}\big(\mathbf{y}(\theta^{\prime}_{i};x),\mathbf{y}(\theta^{\prime}_{j};x)\big)\big]\leq\sigma,$
for $x\in\mathcal{D}^{w}_{i}$ or $x\in\mathcal{D}^{w}_{j}$, where $\mathtt{Div}(\cdot)$ is a divergence measurement function (e.g., KL divergence), and $\sigma$ is a predefined collision threshold.
@@
$\mathbb{E}_{x\in\mathcal{D}^{w}_{i}}\big[\mathtt{Div}\big(\mathbf{y}(\theta^{\prime}_{i};x),\mathbf{y}(\theta^{\prime}_{j};x)\big)\big]>\sigma,\quad\forall j\in[n],j\neq i.$
@@
Problem Formulation. We define the problem of injecting traceable black-box watermarks in FL as in Problem 1.
@@
$s.t.\ \mathbb{E}_{x\in\mathcal{D}^{w}_{i}}$ $[\mathtt{Div}(\mathbf{y}(\theta+\delta_{i};x),\mathbf{y}(\theta+\delta_{j};x))]>\sigma,\quad\forall i,j\in[n],\ i\neq j,$
@@
We derive the following key insights, which motivate the design of our method:
1) A straightforward solution for Problem 1 is to offload each watermark dataset $\mathcal{D}^{w}_{i}$ to client $i$, allowing clients to mix $\mathcal{D}^{w}_{i}$ with their main task dataset $\mathcal{D}^{l}_{i}$ during local training to solve both objectives simultaneously, as proposed in (Nie and Lu, 2024a; b; Liu et al., 2021; Xu et al., 2024; Wu et al., 2022; Li et al., 2022a). However, this method is highly vulnerable to malicious clients who may simply discard $\mathcal{D}^{w}_{i}$, leading to the absence of watermarks in their models. Furthermore, if malicious clients are aware of the watermarking process, they could intentionally tamper with it, undermining its effectiveness. To mitigate these risks, it is preferable to decompose Problem 1, leaving the main task to local clients while performing watermark injection solely on the server.
2) A critical challenge in watermark injection is the risk of watermark collisions due to model averaging during aggregation. Specifically, even if the server successfully injects distinct watermarks into the models before sending them to clients for local training, these watermarks will be fused during model aggregation in the next training round if parameters from all clients are simply averaged, as in FedAvg. To address this issue, a specialized mechanism is required to prevent watermark entanglement during aggregation.
3) The constraint in Problem 1 suggests that to avoid collisions, the server should maximize $\|\delta_{i}-\delta_{j}\|^{2}_{2}$. However, if $\delta_{i}$ and $\delta_{j}$ become too divergent, this may impair the main task performance of the watermarked models. Additionally, since the server lacks access to clients’ local datasets, directly solving the watermarking objective $L(\cdot)$ could also lead to significant degradation in the performance of the main task. Thus, a specialized learning strategy is required to ensure that watermark injection does not compromise the model’s main task performance.
4 Injecting Traceable Black-box Watermarks
Based on the insights in Remark 2, we propose a novel method called $\mathbf{TraMark}$ detailed in Algorithm 1, which can be easily integrated into existing FedAvg frameworks to solve Problem 1. We give the complete process of FedAvg with $\mathbf{TraMark}$ in Algorithm 2 in the Appendix A.3.
4.1 Constraining Watermarking Region